OpenStudy (anonymous):

3squareroot2-2squareroot3 over 3squareroot2 +2squareroot3

OpenStudy (anonymous):

\[3\sqrt{2}-2\sqrt{3}\over3\sqrt{2}+2\sqrt{3}\]